OpenStudy (anonymous):

Estimate the area of a square floor if each side is 12 feet, 9 inches long. Use whole numbers. Enter the numerical answer only.

OpenStudy (anonymous):

12ft 9in is almost 13ft area = s^2 13*13 = 169 12*12 = 144 so it is closer to 169 than 144 , and in between them